# Normal generators for mapping class groups

### v1

@@ Abstract.
In this expository note, we discuss normal generators for mapping class groups of surfaces. Especially, we focus on the relation between normal generation of a mapping class with its asymptotic translation lengths on the Teichmüller space and the curve graph of the underlying surface. We also discuss several open questions.
@@ 1. Introduction
Ever since Thurston brought it to prominence [Thu88], the mapping class group of a surface has become a ubiquitous object in the study of geometry, topology, and dynamics in low-dimensional settings. Formally speaking, given a surface, the mapping class group is defined as the group of orientation preserving homeomorphisms, modulo the subgroup consisting of homeomorphisms isotopic to the identity. For a general background on the subject, see books of Farb and Margalit [FM12] and of Minsky [Min13].
For one thing, the quotient of the Teichmüller space by the mapping class group is the moduli space of algebraic curves. More precisely, let $S_{g}$ be the closed orientable connected surface of genus $g\geq 2$. The Teichmüller space $\mathcal{T}(S_{g})$ is the space of all marked hyperbolic structures on $S_{g}$. Then the mapping class group $\operatorname{Mod}(S_{g})$ acts properly discontinuously on $\mathcal{T}(S_{g})$ by isometries and the quotient is the moduli space $\mathcal{M}_{g}$ of algebraic curves of genus $g$. In fact, since $\mathcal{T}(S_{g})$ is simply connected, the Teichmüller space $\mathcal{T}(S_{g})$ is the universal cover of the moduli space $\mathcal{M}_{g}$ and $\operatorname{Mod}(S_{g})$ is the orbifold fundamental group of $\mathcal{M}_{g}$.
@@
One of the most famous examples of proper normal subgroups of the mapping class group is so-called Torelli group $\mathcal{I}_{g}$. The Torelli group is defined as the subgroup of $\operatorname{Mod}(S_{g})$ of elements which act trivially on $H_{1}(S_{g}):=H_{1}(S_{g};\mathbb{Z})$. It is easy to see that Dehn twists along separating curves are elements of this group (see Example 2.2 for the definition of Dehn twist). Another well-known type of elements in the Torelli group is the so-called bounding pair map. It is of the form $T_{b_{1}}\circ T_{b_{2}}^{-1}$ where $b_{1},b_{2}$ are disjoint homologous simple closed curves and $T_{b_{i}}$ denotes the Dehn twist along the curve $b_{i}$, $i=1,2$. In fact, the Torelli group is generated by Dehn twists along separating curves and bounding pair maps. Another famous example is the Johnson kernel $\mathcal{K}_{g}$. It is the kernel of the Johnson homomorphism and is the subgroup of $\mathcal{I}_{g}$ generated by Dehn twists along separating curves. In fact, there exists a whole filtration of proper normal subgroups including these examples, so-called Johnson filtration but we are not going into details about it in this note.
@@
One can ask a slightly different question, namely, what elements can a proper normal subgroup of $\operatorname{Mod}(S_{g})$ have? In fact, the question we would like to focus on in this note is the opposite question: which elements of $\operatorname{Mod}(S_{g})$ are never contained in any proper normal subgroups? From the point of view of $\operatorname{Mod}(S_{g})=\pi_{1}(\mathcal{M}_{g})$, one can interpret the question as asking which closed curves in the moduli space never lift to a closed curve in any regular cover of the moduli space. In general, an element $h$ of a group $G$ is called a normal generator if its normal closure $\langle\langle h\rangle\rangle$ is the entire group $G$, i.e., no proper normal subgroup contains the given element $h$. Hence, our aim is to find normal generators of the mapping class groups.
Maher and Tiozzo proved that normal generators of mapping class groups are not generic [MT21], by showing that the normal closure of a random mapping class is a free group. On the other hand, surprisingly, it was shown by Lanier and Margalit [LM22] that normal generation of a mapping class turns out to be related to its (asymptotic) translation length on the Teichmüller space: for $f\in\operatorname{Mod}(S_{g})$, its translation length on $\mathcal{T}(S_{g})$ is defined as
@@
where $d_{\mathcal{T}}$ is the Teichmüller distance on $\mathcal{T}(S_{g})$ and $o\in\mathcal{T}(S_{g})$. More precisely, they showed the following criterion for mapping classes to be normal generators: pseudo-Anosov mapping classes with small translation lengths on Teichmüller spaces are normal generators. Together with the work [Pen91] of Penner, it follows that pseudo-Anosov normal generators are abundant in the following sense.
Theorem 1.1 (Lanier-Margalit).
@@
Lanier and Margalit proved the above theorem by giving a sufficient and necessary condition for a given mapping class to be a normal generator, which we will discuss in Section 3. In our work with Wu [BKW21b], we extended their theorems to certain reducible mapping classes. Before presenting the statement, we first note that if $f\in\operatorname{Mod}(S_{g})$ satisfies $\ell_{\mathcal{T}}(f)>0$, then there exists a subsurface $A\subset S_{g}$ invariant under some power of $f$ and its restriction on $A$ is pseudo-Anosov.
Theorem 1.2 (Baik-Kim-Wu).
Let $f\in\operatorname{Mod}(S_{g})$ be such that $f$ preserves a subsurface $A\subseteq S_{g}$ of genus at least three and $f|_{A}$ is pseudo-Anosov. If $\ell_{\mathcal{T}}(f)\leq\log\sqrt{2}$, then $f$ is a normal generator.
Another metric space on which $\operatorname{Mod}(S_{g})$ naturally acts is the curve graph of $S_{g}$. The curve graph $\mathcal{C}(S_{g})$ is defined as the graph whose vertices are isotopy classes of essential simple closed curves on $S_{g}$ and two vertices are connected by an edge if they have disjoint representatives. The curve graph was first introduced by Harvey [Har81]. We equip $\mathcal{C}(S_{g})$ with a simplicial metric $d_{\mathcal{C}}$. For $f\in\operatorname{Mod}(S_{g})$, its (asymptotic) translation length on $\mathcal{C}(S_{g})$ is defined in the same way:
@@
Masur and Minsky showed in [MM99] that $\mathcal{C}(S_{g})$ is Gromov hyperbolic. Moreover, $\ell_{\mathcal{C}}(f)>0$ if and only if $f$ is pseudo-Anosov. Hence, from the point of view of Theorem 1.1, it is natural to ask whether small $\ell_{\mathcal{C}}(\cdot)$ implies normal generation (cf. [BKSW23, Question 1.2]). While this question is widely open, the following theorem shows how small it should be to make the question have an affirmative answer:
Theorem 1.3 (Baik-Kim-Wu).
@@
We note that while both $\mathcal{T}(S_{g})$ and $\mathcal{C}(S_{g})$ are metric spaces on which $\operatorname{Mod}(S_{g})$ naturally act, translation lengths measured on them behave quite differently. For instance, Bader recently showed in [Bad25] that for each $g\geq 2$, there exists a sequence of pseudo-Anosov $f_{n}\in\operatorname{Mod}(S_{g})$ such that
@@
This article is devoted to an exposition of the study of the relation between normal generators of mapping class groups and translation lengths on Teichmüller spaces and curve graphs. In the rest of this article, we discuss some key ideas in the proofs of theorems introduced above. In the last section, we also discuss some further questions, which are widely open.
@@ Acknowledgements
We would like to extend our gratitude to Chenxi Wu for his valuable collaboration with the authors on the primary work discussed in this article. Baik was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) RS-2025-00513595.
@@ 2. Mapping class groups
By a surface, we mean a connected oriented surface of finite genus, possibly with finitely many punctures, and we further assume that its Euler characteristic is negative. The major object of this article is the mapping class group of a surface, and this section is devoted to a brief overview on mapping class groups. We refer to [FM12] and [Min13] for comprehensive references.
@@
One example of a pure mapping class is a Dehn twist. Let $A=\{re^{i\theta}:r\in[1,3],\theta\in[0,2\pi]\}$ be an annulus in the complex plane equipped with the standard orientation on it. Let $h:A\to A$ be an orientation preserving homeomorphism defined as
@@
For a simple closed curve $c$ on $S$, the Dehn twist $T_{c}$ along $c$ is a mapping class defined as follows: let $A_{c}\subset S$ be a neighborhood of $c$ such that there exists an orientation preserving homeomorphism $f_{c}:A_{c}\to A$ that maps $c$ to a circle $\{2e^{i\theta}:\theta\in[0,2\pi]\}$. Then $f_{c}^{-1}\circ h\circ f_{c}:A_{c}\to A_{c}$ is an orientation preserving homeomorphism, whose restriction on $\partial A_{c}$ is the identity. We then extend $f_{c}^{-1}\circ h\circ f_{c}$ to the entire surface $S$, and the mapping class of this extension is the Dehn twist $T_{c}$ along $c$. See Figure 2.
$d$$c$
$d$$c$$c$
$d$$c$$c$
@@
Note that two isotopic simple closed curves have the same Dehn twist. Hence a Dehn twist can be regarded to be along an isotopy class of a simple closed curve. A simple closed curve on $S$ is called essential if it is not homotopic to a point or a puncture. One can see that the Dehn twist $T_{c}$ is non-trivial if and only if $c$ is essential. It is easy to see that for $f\in\operatorname{Mod}(S)$,
@@
Since each Dehn twist $T_{c_{i}}$ is supported in a neighbhorhood of $c_{i}$, it follows from the disjointness among $c_{1},\cdots,c_{k}$ that $T_{c_{1}},\cdots,T_{c_{k}}$ commute. Hence, the above composition is well-defined, meaning that the product does not depend on the labeling of $c_{1},\cdots,c_{k}$.
Dehn twists are not only the simplest mapping classes, but also the fundamentals. It is a classical theorem of Dehn [Deh87] and Lickorish [Lic64] that $\operatorname{PMod}(S)$ is generated by a finitely many Dehn twists. To state the theorem in a more informative way, we introduce some notions. A simple closed curve is called non-separating if its complement is connected, and separating otherwise. For two simple closed curves $c,d$, their geometric intersection number is defined as
@@
There are finitely many non-separating simple closed curves $c_{1},\cdots,c_{k}$ such that $i(c_{i},c_{j})\leq 1$ for all $i,j\in\{1,\cdots k\}$ and Dehn twists $T_{c_{1}},\cdots,T_{c_{k}}$ generate $\operatorname{PMod}(S)$.
Nielsen and Thurston classified mapping classes into three categories ([Nie44], [Thu88]).
Theorem 2.5 (Nielsen-Thurston classification).
@@
    $f$ is pseudo-Anosov, i.e., there exist a representative $f_{0}$ of the isotopy class $f$ and a pair of transverse measured foliations that are invariant under $f_{0}$ and their transverse measures are multiplied by $\lambda_{f}$ and $1/\lambda_{f}$ for some $\lambda_{f}>1$. In this case, the constant $\lambda_{f}$ is called the stretch factor of $f$.
More detailed discussion on measured foliations and transverse measures is required to precisely understand what pseodo-Anosov means. Howeover, it is enough for us to know the classification theorem and the fact that no power of a pseudo-Anosov mapping class fixes an isotopy class of a simple closed curve, in contrast to reducible mapping classes.
@@
As mentioned in the introduction, we are interested in normal subgroups. Given a group $G$, one way to produce its normal subgroup is considering the commutator subgroup of $G$. For $h,k\in G$, we set
$[h,k]:=hkh^{-1}k^{-1}\in G,$
@@
is the commutator subgroup of $G$, and it is easy to see that the commutator subgroup is a normal subgroup. The quotient $G/[G,G]$ is abelian, and is called the abelianization of $G$. The group $G$ is called perfect if
@@
$[\operatorname{PMod}(S),\operatorname{PMod}(S)]=\operatorname{PMod}(S).$
@@
To do this, we introduce the notion of algebraic intersection number. For $c,c^{\prime}\in\pi_{1}(S)$, their algebraic intersection number $\hat{i}(c,c^{\prime})$ is defined as the sum of the indices of the intersection points of $c$ and $c^{\prime}$, where the index of an intersection point is $1$ if the orientation of $c$ and $c^{\prime}$ at the intersection coincides with the orientation of the surface, and $-1$ otherwise. Note that $\hat{i}(c,c^{\prime})$ depends only on homology classes of $c$ and $c^{\prime}$ in $H_{1}(S)$, and hence we use the same notation $\hat{i}(\cdot,\cdot)$ when it is discussed for homology classes.
@@
Note that while the definition of Dehn twist does not involve the choice of orientation on a simple closed curve, we let $c$ to be oriented above. This is only for considering the homology class of $[c]$. Indeed, the above expression does not depend on the orientation on $c$.
@@
Let $S=S_{g,n}$ be a surface of genus $g$ with $n$ number of punctures. The Torelli group $\mathcal{I}_{g,n}$ is the subgroup of $\operatorname{PMod}(S)$ consisting of elements that act trivially on the first homology $H_{1}(S)$.
In other words, the Torelli group $\mathcal{I}_{g,n}$ is the kernel of the canonical homomorphism $\operatorname{PMod}(S)\to\operatorname{Aut}(H_{1}(S))$, and hence is a normal subgroup. An element $f\in\operatorname{PMod}(S)$ is called Torelli, or a Torelli element, if $f\in\mathcal{I}_{g,n}$. Note that the Torelli group is non-trivial and proper subgroup of $\operatorname{PMod}(S)$. Indeed, it follows from (2.1) that for an essential simple closed curce $c$, the Dehn twist $T_{c}$ is Torelli if and only if $c$ is separating. Thurston also showed that there are pseudo-Anosov Torelli elements [Thu88].
@@
In fact, the existence of Torelli pseudo-Anosov mapping class is a consequence of Thurston’s explicit construction of pseudo-Anosov mapping classes [Thu88]. Thurston’s construction of pseudo-Anosov mapping class was generalized by Penner [Pen88], while Thurston’s construction is not restricted to pseudo-Anosovs. We close this section by introducing Penner’s construction, which is more combinatorial.
Let $S$ be a closed surface of genus $g\geq 2$, for convenience. Let $c_{1},\cdots,c_{k}$ be disjoint simple closed curves on $S$ and $c=\cup_{i=1}^{k}c_{i}$ be the multicurve on $S$. Similarly, let $d_{1},\cdots,d_{m}$ be disjoint simple closed curves on $S$ and set $d=\cup_{i=1}^{m}d_{i}$. We say that $c$ and $d$ fill the surface $S$ (or $c_{1},\cdots,c_{k}$ and $d_{1},\cdots,d_{k}$ fill $S$) if every component of $S-(c\cup d)$ is an open disk. In other words, any essential simple closed curve on $S$ must intersect $c\cup d$. See Figure 3.
@@
3. Lanier-Margalit’s well-suited criterion and its generalizations
For a group $G$ and its subset $H\subset G$, we denote by $\langle\langle H\rangle\rangle$ the normal closure of $H$ in $G$, which is the smallest normal subgroup of $G$ containing $H$. An element $h\in G$ is called a normal generator of $G$ if its normal closure $\langle\langle h\rangle\rangle$ is the whole group $G$. Lanier and Margalit [LM22] provided a criterion for a mapping class to be a normal generator of the mapping class group, by considering a certain graph defined for a given mapping class. This criterion is called well-suited criterion.

### v2

@@ Abstract.
In this chapter, we discuss normal generators for mapping class groups of surfaces. Especially, we focus on the relation between normal generation of a mapping class with its asymptotic translation lengths on the Teichmüller space and the curve graph of the underlying surface. We also discuss several open questions.
@@ 1. Introduction
Ever since Thurston brought it to prominence [THU88], the mapping class group of a surface has become a ubiquitous object in the study of geometry, topology, and dynamics in low-dimensional settings. He classified homeomorphisms of surfaces, a classification which divides mapping classes into three types: periodic, reducible, and pseudo-Anosov ([NIE44], [THU88]). This classification, now called Nielsen–Thurston classification, provides a powerful and fundamental tool for understanding the dynamics of surface diffeomorphisms. His work [THU86] also introduced the concept of Thurston norm on the homology of 3-manifolds, which has deep connections to surface theory and mapping class groups.
Formally speaking, given a surface, the mapping class group is defined as the group of orientation preserving homeomorphisms, modulo the subgroup consisting of homeomorphisms isotopic to the identity. For a general background on the subject, see the books of Farb and Margalit [FM12] and of Minsky [MIN13].
For one thing, the quotient of the Teichmüller space by the mapping class group is the moduli space of algebraic curves. More precisely, let $S_{g}$ be the closed orientable connected surface of genus $g\geq 2$. The Teichmüller space $\mathcal{T}(S_{g})$ is the space of all marked hyperbolic structures on $S_{g}$. Then the mapping class group $\operatorname{Mod}(S_{g})$ acts properly discontinuously on $\mathcal{T}(S_{g})$ by isometries and the quotient is the moduli space $\mathcal{M}_{g}$ of algebraic curves of genus $g$. In fact, since $\mathcal{T}(S_{g})$ is simply connected, the Teichmüller space $\mathcal{T}(S_{g})$ is the orbifold universal cover of the moduli space $\mathcal{M}_{g}$ and $\operatorname{Mod}(S_{g})$ is the orbifold fundamental group of $\mathcal{M}_{g}$.
@@
One of the most famous examples of proper normal subgroups of the mapping class group is the so-called Torelli group $\mathcal{I}_{g}$. The Torelli group is defined as the subgroup of $\operatorname{Mod}(S_{g})$ of elements which act trivially on $H_{1}(S_{g}):=H_{1}(S_{g};\mathbb{Z})$. It is easy to see that Dehn twists along separating curves are elements of this group (see Example 2.2 for the definition of Dehn twist). Another well-known type of elements in the Torelli group is the so-called bounding pair map. It is of the form $T_{b_{1}}\circ T_{b_{2}}^{-1}$ where $b_{1},b_{2}$ are disjoint homologous simple closed curves and $T_{b_{i}}$ denotes the Dehn twist along the curve $b_{i}$, $i=1,2$. In fact, the Torelli group is generated by Dehn twists along separating curves and bounding pair maps. Another famous example is the Johnson kernel $\mathcal{K}_{g}$. This is the kernel of the Johnson homomorphism and it is the subgroup of $\mathcal{I}_{g}$ generated by Dehn twists along separating curves. In fact, there exists a whole filtration of proper normal subgroups including these examples, called Johnson filtration, but we are not going into details about it in this chapter.
@@
One can ask a slightly different question, namely, what elements can a proper normal subgroup of $\operatorname{Mod}(S_{g})$ have? In fact, the question we would like to focus on in this chapter is the opposite question: which elements of $\operatorname{Mod}(S_{g})$ are never contained in any proper normal subgroups? From the point of view of $\operatorname{Mod}(S_{g})=\pi_{1}(\mathcal{M}_{g})$, one can interpret this question as asking which closed curves in the moduli space never lift to a closed curve in any regular cover of the moduli space. In general, an element $h$ of a group $G$ is called a normal generator if its normal closure $\langle\langle h\rangle\rangle$ is the entire group $G$, i.e., no proper normal subgroup contains the given element $h$. Hence, our aim is to find normal generators of the mapping class groups.
Maher and Tiozzo proved that normal generators of mapping class groups are not generic [MT21], by showing that the normal closure of a random mapping class is a free group. This suggests that finding a normal generator of $\operatorname{Mod}(S_{g})$ is a challenging task. On the other hand, Lickorish [LIC64] and Mumford [MUM67] showed that the Dehn twist along a non-separating simple closed curve on $S_{g}$, which is obtained by cutting, rotating, and regluing the surface along the curve, is a normal generator of $\operatorname{Mod}(S_{g})$. Many periodic normal generators are also given by Harvey and Korkmaz ([HK05], [KOR05]), Lanier [LAN18], McCarthy and Papadopoulos [MP87], and Yoshihara [YOS16]. Indeed, Lanier and Margalit [LM22] proved that every non-trivial periodic mapping class that is not a hyperelliptic involution is a normal generator of $\operatorname{Mod}(S_{g})$ when $g\geq 3$.
For pseudo-Anosov mapping classes, surprisingly, it was shown by Lanier and Margalit [LM22] that the normal generation of a mapping class turns out to be related to its (asymptotic) translation length on Teichmüller space: for $f\in\operatorname{Mod}(S_{g})$, its translation length on $\mathcal{T}(S_{g})$ is defined as
@@
where $d_{\mathcal{T}}$ is the Teichmüller distance on $\mathcal{T}(S_{g})$ and $o$ is an arbitrary point in $\mathcal{T}(S_{g})$. More precisely, they showed the following criterion for mapping classes to be normal generators: pseudo-Anosov mapping classes with small translation lengths on Teichmüller spaces are normal generators. Together with the work [PEN91] of Penner, it follows that pseudo-Anosov normal generators are abundant in the following sense.
Theorem 1.1 (Lanier–Margalit).
@@
Lanier and Margalit proved this theorem by giving a sufficient and necessary condition for a given mapping class to be a normal generator, which we will discuss in Section 3. In our work with Wu [BKW25], we extended their theorems to certain reducible mapping classes. Before presenting the statement, we first note that if $f\in\operatorname{Mod}(S_{g})$ satisfies $\ell_{\mathcal{T}}(f)>0$, then there exists a subsurface $A\subset S_{g}$ which is invariant under some power of $f$ and whose restriction to $A$ is pseudo-Anosov.
Theorem 1.2 (Baik–Kim–Wu).
Let $f\in\operatorname{Mod}(S_{g})$ be a mapping class that preserves a subsurface $A\subseteq S_{g}$ of genus at least three and suppose that $f|_{A}$ is pseudo-Anosov. If $\ell_{\mathcal{T}}(f)\leq\log\sqrt{2}$, then $f$ is a normal generator.
Another metric space on which $\operatorname{Mod}(S_{g})$ naturally acts is the curve graph of $S_{g}$. This graph $\mathcal{C}(S_{g})$ is defined as the graph whose vertices are isotopy classes of essential simple closed curves on $S_{g}$ and where two vertices are connected by an edge if they have disjoint representatives. The curve graph was first introduced by Harvey [HAR81]. We equip $\mathcal{C}(S_{g})$ with a simplicial metric $d_{\mathcal{C}}$. For $f\in\operatorname{Mod}(S_{g})$, its (asymptotic) translation length on $\mathcal{C}(S_{g})$ is defined in the same way:
@@
Masur and Minsky showed in [MM99] that $\mathcal{C}(S_{g})$ is Gromov hyperbolic. Moreover, $\ell_{\mathcal{C}}(f)>0$ if and only if $f$ is pseudo-Anosov. Hence, from the point of view of Theorem 1.1, it is natural to ask whether small $\ell_{\mathcal{C}}(\cdot)$ implies normal generation (cf. [BKS+23, Question 1.2]). While this question is wide open, the following theorem shows how small it should be to make the question have an affirmative answer:
Theorem 1.3 (Baik–Kim–Wu).
@@
We note that while both $\mathcal{T}(S_{g})$ and $\mathcal{C}(S_{g})$ are metric spaces on which $\operatorname{Mod}(S_{g})$ naturally acts, translation lengths measured on them behave quite differently. For instance, Bader recently showed in [BAD25] that for each $g\geq 2$, there exists a sequence of pseudo-Anosovs $f_{n}\in\operatorname{Mod}(S_{g})$ such that
@@
This chapter is devoted to the study of the relation between normal generators of mapping class groups and translation lengths on Teichmüller spaces and curve graphs. In the rest of this chapter, we discuss some key ideas in the proofs of the theorems introduced above. In the last section, we also discuss some further questions which are wide open.
Organization
In Section 2, we review structural aspects of mapping class groups. Section 3 is devoted Lanier–Margalit’s well-suited criterion and its generalization, which is crucial in the study of normal generation of mapping class groups. We discuss the relation between translation lengths of Teichmüller spaces and normal generation in Section 4. In Section 5, we move on to the discussion of asymptotic translation lengths on curve graphs. We record various further questions in Section 6.
@@ Acknowledgements
We would like to extend our gratitude to Chenxi Wu for his valuable collaboration with the authors on the primary work discussed in this chapter. Baik was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) RS-2025-00513595.
@@ 2. Mapping class groups
By a surface, we mean a connected oriented surface of finite genus, possibly with finitely many punctures, and we further assume that its Euler characteristic is negative. The major object of this chapter is the mapping class group of a surface, and this section is devoted to a brief overview of mapping class groups. We refer to [FM12] and [MIN13] for comprehensive references.
@@
One example of a pure mapping class is a Dehn twist. Intuitively, a Dehn twist is constructed by cutting the surface along a simple closed curve, rotating the surface along the curve, and then regluing it. To be precise, let $A=\{re^{i\theta}:r\in[1,3],\theta\in[0,2\pi]\}$ be an annulus in the complex plane equipped with the standard orientation on it. Let $h:A\to A$ be an orientation preserving homeomorphism defined as
@@
For a simple closed curve $c$ on $S$, the Dehn twist $T_{c}$ along $c$ is a mapping class defined as follows: let $A_{c}\subset S$ be a neighborhood of $c$ equipped with an orientation-preserving homeomorphism $f_{c}:A_{c}\to A$ that maps $c$ to a circle $\{2e^{i\theta}:\theta\in[0,2\pi]\}$. Then $f_{c}^{-1}\circ h\circ f_{c}:A_{c}\to A_{c}$ is an orientation preserving homeomorphism whose restriction on $\partial A_{c}$ is the identity. We then extend $f_{c}^{-1}\circ h\circ f_{c}$ to the entire surface $S$. The mapping class of this extension is the Dehn twist $T_{c}$ along $c$. See Figure 2. As in Theorem 2.4 below, Dehn twists form a finite generating set of $\operatorname{PMod}(S)$.
ddcc
ddcccc
ddcccc
@@
Note that two isotopic simple closed curves define the same Dehn twist. Hence a Dehn twist can be regarded to be along an isotopy class of a simple closed curve. A simple closed curve on $S$ is called essential if it is not homotopic to a point or a puncture. One can see that the Dehn twist $T_{c}$ is non-trivial if and only if $c$ is essential. It is easy to see that for $f\in\operatorname{Mod}(S)$, we have
@@
Since each Dehn twist $T_{c_{i}}$ is supported in a neighborhood of $c_{i}$, it follows from the disjointness of $c_{1},\cdots,c_{k}$ that $T_{c_{1}},\cdots,T_{c_{k}}$ commute. Hence, the above composition is well-defined, meaning that the product does not depend on the labeling of $c_{1},\cdots,c_{k}$.
Dehn twists are not only the simplest mapping classes, but also the fundamental. It is a classical theorem of Dehn [DEH87] and Lickorish [LIC64] that $\operatorname{PMod}(S)$ is generated by finitely many Dehn twists. To state the theorem in a more informative way, we introduce some notions. An essential simple closed curve is called non-separating if its complement is connected, and separating otherwise. For two simple closed curves $c,d$, their geometric intersection number is defined as
@@
There exist finitely many non-separating simple closed curves $c_{1},\cdots,c_{k}$ such that $i(c_{i},c_{j})\leq 1$ for all $i,j\in\{1,\cdots k\}$ and Dehn twists $T_{c_{1}},\cdots,T_{c_{k}}$ generate $\operatorname{PMod}(S)$.
We note that while Dehn and Lickorish had different approaches, the statement of finite generation by Dehn twists along non-separating simple closed curves is referred to as the Dehn–Lickorish theorem. The precise picture of curves $c_{1},\cdots,c_{k}$ satisfying Theorem 2.4 can be found in [LIC64, Fig. 1].
Thurston classified mapping classes into three categories [THU88].
Theorem 2.5 (Thurston’s classification).
@@
    $f$ is pseudo-Anosov, i.e., there exist a representative $f_{0}$ of the isotopy class $f$ and a pair of transverse measured foliations that are invariant under $f_{0}$ and their transverse measures are multiplied by $\lambda_{f}$ and $1/\lambda_{f}$ respectively for some $\lambda_{f}>1$. In this case, the constant $\lambda_{f}$ is called the stretch factor of $f$.
A more detailed discussion on measured foliations and transverse measures is required to precisely understand what pseudo-Anosov means. However, it is enough for us to know the classification theorem and the fact that no power of a pseudo-Anosov mapping class fixes an isotopy class of a simple closed curve, in contrast to reducible mapping classes.
@@
As mentioned in the introduction, we are interested in normal subgroups. Given a group $G$, one way to produce its normal subgroup is to consider the commutator subgroup of $G$. For $h,k\in G$, we set
$h,k]:=hkh^{-1}k^{-1}\in G,$
@@
is the commutator subgroup of $G$. It is easy to see that the commutator subgroup is a normal subgroup. The quotient $G/[G,G]$ is abelian and is called the abelianization of $G$. The group $G$ is called perfect if
@@
$\operatorname{PMod}(S),\operatorname{PMod}(S)]=\operatorname{PMod}(S).$
We also note a well-known fact that $\operatorname{Mod}(S)$ has a trivial center if $S$ is neither a closed surface of genus two nor a torus with one or two punctures. When $S$ is one of those exceptional surfaces, the center of $\operatorname{Mod}(S)$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$, consisting of a hyperelliptic involution [FM12, Section 3.4].
@@
[…]
@@
Note that while the definition of Dehn twist does not involve the choice of an orientation on a simple closed curve, we let $c$ be oriented in this discussion. This is only for considering the homology class of $[c]$. Indeed, the above expression does not depend on the orientation on $c$.
@@
Let $S=S_{g,n}$ be a surface of genus $g$ with $n$ punctures. The Torelli group $\mathcal{I}_{g,n}$ is the subgroup of $\operatorname{PMod}(S)$ consisting of elements that act trivially on the first homology $H_{1}(S)$.
In other words, the Torelli group $\mathcal{I}_{g,n}$ is the kernel of the canonical homomorphism $\operatorname{PMod}(S)\to\operatorname{Aut}(H_{1}(S))$, hence it is a normal subgroup. An element $f\in\operatorname{PMod}(S)$ is called Torelli, or a Torelli element, if $f\in\mathcal{I}_{g,n}$. Note that the Torelli group is non-trivial and a proper subgroup of $\operatorname{PMod}(S)$. Indeed, it follows from (2.1) that for an essential simple closed curve $c$, the Dehn twist $T_{c}$ is Torelli if and only if $c$ is separating. Thurston also showed that there are pseudo-Anosov Torelli elements [THU88].
@@
In fact, the existence of Torelli pseudo-Anosov mapping class is a consequence of Thurston’s explicit construction of pseudo-Anosov mapping classes [THU88]. Thurston’s construction of pseudo-Anosov mapping classes was generalized by Penner [PEN88]. We close this section by introducing Penner’s construction, which is more combinatorial.
Let $S$ be a closed surface of genus $g\geq 2$, for convenience. Let $c_{1},\cdots,c_{k}$ be disjoint simple closed curves on $S$ and $c=\cup_{i=1}^{k}c_{i}$ be a multicurve on $S$. Similarly, let $d_{1},\cdots,d_{m}$ be disjoint simple closed curves on $S$ and set $d=\cup_{i=1}^{m}d_{i}$. We say that $c$ and $d$ fill the surface $S$ (or $c_{1},\cdots,c_{k}$ and $d_{1},\cdots,d_{k}$ fill $S$) if every component of $S-(c\cup d)$ is an open disk, when $c$ and $d$ are in minimal position with respect to intersection number. In other words, any essential simple closed curve on $S$ must intersect $c\cup d$. See Figure 3.
@@
3. Lanier–Margalit’s well-suited criterion and its generalizations
For a group $G$ and a subset $H\subset G$, we denote by $\langle\langle H\rangle\rangle$ the normal closure of $H$ in $G$. This is the smallest normal subgroup of $G$ containing $H$. An element $h\in G$ is called a normal generator of $G$ if its normal closure $\langle\langle h\rangle\rangle$ is the whole group $G$. Lanier and Margalit [LM22] provided a criterion for a mapping class to be a normal generator of the mapping class group, by considering a certain graph defined for a given mapping class. This criterion is called well-suited criterion.